A particular brand of tires claims that its deluxe tire averages at least 50,000 miles before it needs to be replaced. From past studies of this tire, the standard deviation is known to be 8,000. A survey of owners of that tire design is conducted. Of the 32 tires surveyed, the mean lifespan was 46,600 miles with a standard deviation of 9,800 miles. Using alpha = 0.05, is the data highly consistent with the claim?

Note: If you are using a Student's t-distribution for the problem, you may assume that the underlying population is normally distributed. (In general, you must first prove that assumption, though.)

O Part (d) State the distribution to use for the test. (Round your answers to two decimal places.) Х 50000 1511.86 n 1x ) 1 PPart (h) Indicate the correct decision (reject or do not reject the null hypothesis), the reason for it, and write an appPart (0) Construct a 95% confidence interval for the true mean. Sketch the graph of the situation. Label the point estimate a
Answer #1

Solution : given, population mean M= 50000 population Standard deviation, 6 = 8000 Samble Size ne 32 Sample mean, 4 6600 SampHot M > 5000 o (claim) Hi : M < 50000 (left tailed A les native) This is one failed test. Under Ho, test statistic Z- 46600 5Past i Confidence level = 95% = 0.95 Significance level = 1-0.95 = 0.05 95% confidence interval for true mean (M) M х + Zo Nъu 44 600 + 1.46 xBooo 22 - - u M ч 66 oo + 1.46 x 1ч1 ч. 21) (46600 211). 85 ) 46 600 — 21+. 85, 966 oot 217 • 85 цE (458 28.